Hello could someone please help me with this equation, I am so stuck! I have tried to make a formula but it is not working. I tried (x*.2)+(y*.5)=25.30 but this formula does not include the 68 because I do not know where to put it. The equation is below.
Mike saves all his 50 cent & 20 cent coins. When he counts them he finds he has 68 coins amounting to $25.30. How many of each type of coin does he have?

Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

Mike saves all his 50 cent & 20 cent coins. When he counts them he finds he has 68 coins amounting to $25.30. How many of each type of coin does he have?
------------
f + t = 68
50f + 20t = 2530
=======

Answer by mananth(16946)

x 20 cent coins
y 50 cent coins
Total attended
1 x + 1 y = 68.00 .............1

0.20 x 0.50 y = 25.30 .............2
Eliminate y
multiply (1)by -0.50
Multiply (2) by 1.00
-0.50 x -0.50 y = -34.00
0.20 x + 0.50 y = 25.30
Add the two equations
-0.30 x = -8.70
/ -0.30
x = 29.00
plug value of x in (1)
1.00 x + 1.00 y = 68.00
29.00 + 1.00 y = 68.00
1.00 y = 68.00 -29.00
1.00 y = 39.00
y = 39.00
29 20 cent coins
39 50 cent coins
m.ananth@hotmail.ca